two cars going in opposite direction. car B travel 25mph faster then car A. in 6 hours the cars are 620 mile apart. find the speed of each car.
Answer by nerdybill(7384)

two cars going in opposite direction. car B travel 25mph faster then car A. in 6 hours the cars are 620 mile apart. find the speed of each car.
.
Let x = speed of slower car
then
x+25 = speed of faster car
.
applying the distance formula:
6(x + x+25) = 620
6(2x+25) = 620
12x+150 = 620
12x = 470
x 39.17 mph (slower car)
.
faster car:
x+25 = 39.17+25 = 64.17 mph